e posnts . A.3e g sampie or a nyarocaroon compouno containing onay C ana nyarogen pr of CO2. A 9.29 ml sample weighing 21.6 ㎎was colected over water at a pressure of 796 torr at 25°C nyarogen proaucea cu.s (The vapor pressure of water is 23.8 torr at 25C) What is the molecular formula of the compound? (8 points) 5. The molar heat capacity of Ag is 25.35 J/mol C. What is the final temperature of a 123.5 g piece of silver if the initial temperature was 25.5 °C and 31 KJ of heat was added? (6 points) 6. Is the following, Polar, nonpolar, or ionic compounds? a. O-Si-O Page 5 Need help with questions 4, 5,6. Can you explain how to and why we do certain steps to solve.
Answer #1

1st question is not quite visible. starting part.Not sure whether it is 6.52 or 6.92

I will answer the rest for you

5)
Q = 31 KJ = 31000 J
mass of silver = 123.5 g
number of moles of silver, n =mass/molar mass
= 123.5 / 107.87
= 1.145 mol

use:
Q = n*C*delta T
31000 = 1.145*25.35*(T - 25.5)
(T - 25.5) = 1068
T = 1093.5 oC
Answer: 1093.5 oC

6)
It is covalent and non polar
Due to symmetry on both sides, net dipole moment is zero.
This is what makes it non polar